# Stochastic Thermodynamics of Non-reciprocally Interacting Particles and Fields

### v2

@@ D Temporal cross-correlations between macrostates and relaxation EP
We define the temporal correlation $C_{ij}(t,\Delta t)=\rho_{i}(t+\Delta t)\rho_{j}(t)$ between the macrostates and its anti-symmetric $C_{ij}^{a}(t,\Delta t)=C_{ji}(t,\Delta t)-C_{ij}(t,\Delta t)$ and symmetric $C_{ij}^{s}(t,\Delta t)=C_{ji}(t,\Delta t)+C_{ij}(t,\Delta t)$ decomposition [158]. It can be trivially verified $C_{ij}^{a}(t,\Delta t)=\Delta t\>\omega_{ij}$ for small $\Delta t$. Integrating eq. 37 from an initial time $t_{i}$ to time $t_{f}$ with $\tau=t_{f}-t_{i}$, leads to the following expression for the relaxation process:
@@
eq. 42 relates EP due to reciprocal $\langle E^{int}\rangle$ and nonreciprocal $\langle\dot{\Sigma}^{nr}\rangle$ interactions between the macrostates with symmetric and anti-symmetric temporal correlations between the initial and final state. $\langle{C}_{ij}^{s}(t_{i},\tau)\rangle,\langle{C}_{ij}^{a}(t_{i},\tau)\rangle$ are convenient to obtain experimentally. Using eq. 42 in eq. 40, one obtains a tighter bound on $\langle\dot{\Sigma}\rangle$ by using the relaxation of ${C}_{ij}^{s}(t_{i},\tau)$ and ${C}_{ij}^{a}(t_{i},\tau)$ from the initial state to the final state. This bound is similar to TUR but is obtained using macrostate correlations instead of the precision of the transition currents [159]. Hence, it should be compared to Ref.[160, 161, 162, 163, 164, 165, 166]. $-\langle\dot{E}^{int}\rangle>0$ ensures that attractive (repulsive) reciprocal interactions $V_{ij}^{r}<0$ ($V_{ij}^{r}>0$) increase (decrease) the symmetric temporal macrostate correlation $\langle{C}_{ij}^{s}(t_{i},\tau)\rangle$ for the relaxation process. Similarly, $\langle\dot{\Sigma}^{nr}\rangle>0$ implies that $\langle{{C}_{ij}^{a}(t_{i},\tau)\rangle}\rangle$ increases for $V_{ij}^{nr}>0$.
Importantly, this highlights the physical correctness of ‘the orthogonal gauge’. The choice of any other gauge will incorrectly quantify a part of the symmetric macrostate correlations in $\Sigma^{nr}$, which is physically contradictory. Subsequently, this implies that $E^{int}$ and $\Sigma^{nr}$ are not linearly independent for other gauges. Thus, a subsequent redefinition of the linearly independent contributions of the EPR leads to a formulation equivalent to ‘the orthogonal gauge’ fixing.
@@ 3.4 Mesoscopic, Macroscopic and Deterministic limits
The parameter $\Omega$ dictates the scale of the coarse-grained description of the system. It consists of three important limits, namely mesoscopic, macroscopic and deterministic, which are characterized respectively by $\Omega$ being one, large and infinite. Importantly, the intensiveness of the microscopic Boltzmann weight $\epsilon_{i}^{\#}$ requires the scaling constraint on $v_{ij}^{r},v_{ij}^{nr}\propto 1/\Omega$. The Taylor series expansion of eq. 25 in $v_{ij}^{r},v_{ij}^{nr}$ (or equivalently, in $1/\Omega$) leads to the interaction coefficients in the macroscopic limit, $V_{ij}^{r}=\beta\Omega v_{ij}^{r}+\frac{1}{2}\Omega\beta^{2}\left((v_{ij}^{r})^{2}+(v_{ij}^{nr})^{2}\right)+O(\Omega\beta^{3}v_{ij}^{3})$ and $V_{ij}^{nr}=\beta\Omega v_{ij}^{nr}+\Omega\beta^{2}v_{ij}^{nr}v_{ij}^{r}+O(\Omega\beta^{3}v_{ij}^{3})$. The macroscopic interaction coefficients satisfy the gauge-fixing $V_{ij}^{r}=V_{ji}^{r}$ and $V_{ij}^{nr}=-V_{ji}^{nr}$. Taking $\Omega\to\infty$ gives the deterministic limit, $\bar{V}_{ij}^{r}=\lim_{\Omega\to\infty}\beta\Omega v_{ij}^{r}$ and $\bar{V}_{ij}^{nr}=\lim_{\Omega\to\infty}\beta\Omega v_{ij}^{nr}$. Throughout the paper, we discuss the macroscopic coarse-grained description. However, we should impose the system-specific scale $\Omega$ for the coarse-grained description. In particular, when the mean number of particles per lattice site is small enough, the mesoscopic description is a better alternative to implementing coarse-grained physical analysis.
The table 2 summarizes the implications of fluctuations on observation scales, mesoscopic, macroscopic, and deterministic. The relevant coarse-grained states are the number of particles at a lattice site, the particle density at a lattice point, and the particle density at a lattice point. The suitable physical models where these observation scales are utilized are; the Lattice Gas Models (LGM), Macroscopic Fluctuation Theory (MFT), and Chemical Reaction Networks (CRN), respectively. CRN are the special case where $\Omega=\mathcal{V}$. These structurally similar descriptions vary hugely in other physical aspects, particularly the noise effect’s nature.
Noise plays a key role through two different mechanisms, particle occupancy and the transitions between them. The eq. 25 highlights that the Poissonian occupancy noise renormalizes mesoscopic $V_{ij}$ nonlinearly (in $v_{ij}$). It plays a key role in predicting the microscopic phase diagram using the mesoscopic coarse-grained EOM [86]. Moreover, the mean EPR correctly incorporates the microscopic noise effects of occupancy. ()($\dagger$)( † ) In comparison, the Gaussian/Langevin approximation eq. 28 of the mesoscopic Poissonian transition noise is sufficient due to the van Kampen closure approximation and the correct identification of the fluctuation-response relation discussed subsequently in section 4.2. The non-quadratic dissipation function leads to the non-quadratic Hamilton-Jacobi equation [167, 168, 169, 170, 171]. Hence, eq. 28 incorporates this mesoscopic effect despite the Gaussian/Langevin approximation formulation. A more systematic analysis of mesoscopic Poissonian transition fluctuations is detailed in [88]. Importantly, we have not obtained the macroscopic EPR using the Langevin equation, which avoids the close-to-equilibrium underestimation of the macroscopic EPR using the Langevin equation [88]. In particular, the correct identification of effective transition affinities using the mean and fluctuations of the transition current restores this issue [88].
@@
Occupancy noise Poissonian eq. 25 Gaussian corrections around the mean field Vanishes, recovering the mean-field limit
Transition noise Poissonian and $O(1)$ Gaussian and $O(1/\Omega)$ Vanishes
Table 2: The summary of the impact and importance of the fluctuations across different coarse-grained descriptions.
@@ 4 Thermodynamic relations
Here, we outline different thermodynamic relations using the macroscopic description. However, one could use the microscopic description, which is more relatable in Stochastic Thermodynamics. We deliberately utilize the macroscopic description to extend and exhibit the applicability of Stochastic Thermodynamics to interacting many-body systems.
@@ 4.1 Non-reciprocal Onsager relations
In this section, we demonstrate the non-reciprocal Onsager’s relation [172, 173]. Assuming the vanishing external driving, $\vec{F}_{i}^{sp}=0$ and $F_{\gamma\gamma^{\prime}}^{ch}=0$. The $J^{\vec{\mathcal{D}}}_{i}$ in eq. 29:
$\begin{split}\>\vec{J}^{\vec{\mathcal{D}}}_{i}&=-\sum_{j}\left(D_{ij}^{r}\left(\{\rho\}\right)+D_{ij}^{nr}\left(\{\rho\}\right)\right)\nabla^{\vec{\mathcal{D}}}\rho_{j},\end{split}$ (43)
@@
where, $D_{ij}^{r}(\{\rho\})=D_{i}^{\vec{{\mathcal{D}}}}\left(\{\rho\}\right)(\beta V_{ij}^{r}+\delta_{ij}/\rho_{i})$ and $D_{ij}^{nr}\left(\{\rho\}\right)=\beta D_{i}\left(\{\rho\}\right)V_{ij}^{nr}$. The $D_{ij}^{r}(\{\rho\})$ term satisfies Onsager’s reciprocal relationship [172, 173], as $V_{ij}^{r}={\partial^{2}E}/{\partial\rho_{j}\partial\rho_{i}}={\partial^{2}E}/{\partial\rho_{i}\partial\rho_{j}}$. In contrast, the term $D_{ij}^{nr}(\{\rho\})$ satisfies the Onsager anti-reciprocal relationship, attributed to $V_{ij}^{nr}=-V_{ji}^{nr}$ or $\partial F_{i}^{nr}/\partial\rho_{j}=-\partial F_{j}^{nr}/\partial\rho_{i}$. We introduce the mobility, entropic, symmetric, and anti-symmetric interaction matrices $\mathbb{D}$, $\mathbb{S}$, $\mathbb{D}^{r}$ and $\mathbb{D}^{nr}$, with $ij^{th}$ element of the matrices being $\delta_{ij}D_{i}\left(\{\rho\}\right)$, $\delta_{ij}/\rho_{i}$, $D_{ij}^{r}$ and $D_{ij}^{nr}$, respectively. The column vector $\boldsymbol{\nabla\rho}$.
Thus, $\mathbf{J}^{\vec{\mathcal{D}}}=(\mathbb{D}^{r}+\mathbb{D}^{nr})\nabla\boldsymbol{\rho}=\mathbb{D}(\mathbb{S}+\beta(\mathbb{V}^{r}+\mathbb{V}^{nr}))\nabla\boldsymbol{\rho}$ with the corresponding $\mathbf{F}^{\vec{\mathcal{D}}}=(\mathbb{F}^{r}+\mathbb{F}^{nr})\nabla\boldsymbol{\rho}=\mathbb{D}^{-1}(\mathbb{D}^{r}+\mathbb{D}^{nr})\nabla\boldsymbol{\rho}$. Hence, $\langle\dot{\Sigma}^{\mathcal{D}}\rangle=\langle\mathbf{F}^{\vec{\mathcal{D}}}\cdot\mathbf{J}^{\vec{\mathcal{D}}}\rangle=\boldsymbol{\nabla\rho}^{T}\>(\mathbb{D}^{r}+\mathbb{D}^{nr})^{T}\mathbb{D}^{-1}(\mathbb{D}^{r}+\mathbb{D}^{nr})\boldsymbol{\nabla\rho}$. The square root of $\mathbb{D}^{-1}$ is also a diagonal matrix ${\mathbb{D}^{-\frac{1}{2}}}$ and satisfies the relation with its transpose ${\mathbb{D}^{-\frac{1}{2}}}={\mathbb{D}^{-\frac{1}{2}}}^{T}$. This reduces $\langle\dot{\Sigma}^{\mathcal{D}}\rangle=||\mathbb{D}^{r}+\mathbb{D}^{nr}||^{2}$ to the norm obtained using $\mathbb{D}^{-1}$. In addition, the symmetric and skew-symmetric matrices satisfy ${\mathbb{D}^{r}}^{T}=\mathbb{D}^{r}$ and ${\mathbb{D}^{nr}}^{T}=-\mathbb{D}^{nr}$. Using this $\langle\dot{\Sigma}^{\mathcal{D}}\rangle$ leads to vanishing EPR due to the cross-coupling between the reciprocal and non-reciprocal parts, $\langle\dot{\Sigma}^{\mathcal{D}}\rangle=||\mathbb{D}^{r}||^{2}+||\mathbb{D}^{nr}||^{2}$. Thus, $\langle\dot{\Sigma}^{\mathcal{D}}\rangle=\langle\nabla\boldsymbol{\rho}^{T}({\mathbb{D}^{r}}^{T}\mathbb{D}^{-1}\mathbb{D}^{r}+{\mathbb{D}^{nr}}^{T}\mathbb{D}^{-1}\mathbb{D}^{nr})\nabla\boldsymbol{\rho}\rangle$ or, equivalently $\langle\dot{\Sigma}^{\mathcal{D}}\rangle=\langle\nabla\boldsymbol{\rho}^{T}\>({\mathbb{F}^{r}}^{T}\mathbb{D}^{T}\mathbb{F}^{r}+{\mathbb{F}^{nr}}^{T}\mathbb{D}^{T}\mathbb{F}^{nr})\nabla\boldsymbol{\rho}\rangle$. The $\nabla\boldsymbol{\rho}^{T}{\mathbb{D}^{r}}^{T}\mathbb{D}^{-1}\mathbb{D}^{r}\nabla\boldsymbol{\rho}$ term relates the symmetric response coefficients of $\mathbf{J}^{\vec{\mathcal{D}}}$ to the mean EPR, an analogous formulation of the Onsager’s reciprocal relation close to the equilibrium [172, 173, 174]. The $\nabla\boldsymbol{\rho}^{T}{\mathbb{D}^{nr}}^{T}\mathbb{D}^{-1}\mathbb{D}^{nr}\nabla\boldsymbol{\rho}$ term relates the anti-symmetric response coefficients of $\mathbf{J}^{\vec{\mathcal{D}}}$ to the mean EPR. This is novel formulation of the Onsager non-reciprocal relation. By construction, the norm is positive, leading to two independently positive terms to $\langle\dot{\Sigma}^{\mathcal{D}}\rangle$. Therefore, this decomposition is interpreted as an equivalent representation of the orthogonal decomposition of $\langle\dot{\Sigma}^{\mathcal{D}}\rangle$ for non-reciprocal systems.
@@
Using $\boldsymbol{\rho}=\boldsymbol{\rho}^{ss}+\delta\boldsymbol{\rho}$ or/and $\mathbf{J}^{\vec{\mathcal{D}}}=\mathbf{J}^{\vec{\mathcal{D}}}_{ss}+\delta\mathbf{J}^{\vec{\mathcal{D}}}$, and using the symmetry of the fluctuations around the steady-state profile $\boldsymbol{\rho}^{ss}$ or/and $\mathbf{J}^{\vec{\mathcal{D}}}_{ss}$, one decomposes $\langle\dot{\Sigma}^{\mathcal{D}}\rangle$ further into the housekeeping and nonadiabatic contribution. Similarly, one can trivially incorporate $\langle\dot{\Sigma}^{\mathcal{R}}\rangle$ by linearizing the reactive currents around equilibrium $\boldsymbol{\rho}^{eq}$ ($\mathbf{J}^{\vec{\mathcal{R}}}_{eq}$) or steady-state $\boldsymbol{\rho}^{ss}$ ($\mathbf{J}^{\vec{\mathcal{R}}}_{ss}$), it requires identifying the analogous symmetric and anti-symmetric couplings $\mathbb{D}^{r}$ and $\mathbb{D}^{nr}$ on the discrete transition space [174].

### v3

@@ E Temporal cross-correlations between macrostates and relaxation EP
We define the temporal correlation $C_{ij}(t,\Delta t)=\rho_{i}(t+\Delta t)\rho_{j}(t)$ between macrostates and their anti-symmetric $C_{ij}^{a}(t,\Delta t)=C_{ji}(t,\Delta t)-C_{ij}(t,\Delta t)$ and symmetric $C_{ij}^{s}(t,\Delta t)=C_{ji}(t,\Delta t)+C_{ij}(t,\Delta t)$ decomposition [164]. It can be trivially verified that $C_{ij}^{a}(t,\Delta t)=\Delta t\>\omega_{ij}$ for small $\Delta t$. Integrating eq. 38 from an initial time $t_{i}$ to a final time $t_{f}$, with $\tau=t_{f}-t_{i}$, leads to the following expression for the relaxation process:
@@
Equation 43 relates the EPR due to reciprocal $\langle E^{int}\rangle$ and non-reciprocal $\langle\dot{\Sigma}^{nr}\rangle$ interactions between the macrostates with the symmetric and anti-symmetric temporal correlations between the initial and final states. $\langle{C}_{ij}^{s}(t_{i},\tau)\rangle$ and $\langle{C}_{ij}^{a}(t_{i},\tau)\rangle$ are convenient to obtain experimentally. Using eq. 43 in eq. 41 yields a tighter bound on $\langle\dot{\Sigma}\rangle$ by using the relaxation of ${C}_{ij}^{s}(t_{i},\tau)$ and ${C}_{ij}^{a}(t_{i},\tau)$ from the initial state to the final state. This bound is similar to the TUR but is obtained using macrostate correlations instead of the precision of the transition currents [165]. Hence, it should be compared to Refs. [166, 167, 168, 169, 170, 171, 172]. $\langle\dot{E}^{int}\rangle>0$ ensures that attractive (repulsive) reciprocal interactions $V_{ij}^{r}<0$ ($V_{ij}^{r}>0$) increase (decrease) the symmetric temporal macrostate correlation $\langle{C}_{ij}^{s}(t_{i},\tau)\rangle$ during the relaxation process. Similarly, $\langle\dot{\Sigma}^{nr}\rangle>0$ implies that $\langle{{C}_{ij}^{a}(t_{i},\tau)\rangle}\rangle$ increases for $V_{ij}^{nr}>0$.
Importantly, this highlights the physical correctness of ‘the orthogonal gauge’. The choice of any other gauge would incorrectly assign a part of the symmetric macrostate correlations to $\Sigma^{nr}$, which is physically contradictory. Consequently, $E^{int}$ and $\Sigma^{nr}$ are not linearly independent for other gauges, and a subsequent redefinition of the linearly independent contributions to the EPR leads to a formulation equivalent to the ‘orthogonal gauge’ fixing.
@@ 3.4 Mesoscopic, Macroscopic and Deterministic limits
The parameter $\Omega$ dictates the scale of the coarse-grained description of the system. It consists of three important limits: the mesoscopic, macroscopic, and deterministic limits, characterized respectively by $\Omega$ being one, large, and infinite. Importantly, the intensive scaling of the microscopic Boltzmann weight $\epsilon_{i}^{\#}$ requires the constraint $v_{ij}^{r},v_{ij}^{nr}\propto 1/\Omega$. A Taylor series expansion of eq. 25 in $v_{ij}^{r},v_{ij}^{nr}$ (or equivalently in $1/\Omega$) leads to the macroscopic interaction coefficients $V_{ij}^{r}=\beta\Omega v_{ij}^{r}+\frac{1}{2}\Omega\beta^{2}((v_{ij}^{r})^{2}+(v_{ij}^{nr})^{2})+O(\Omega\beta^{3}v_{ij}^{3})$ and $V_{ij}^{nr}=\beta\Omega v_{ij}^{nr}+\Omega\beta^{2}v_{ij}^{nr}v_{ij}^{r}+O(\Omega\beta^{3}v_{ij}^{3})$. The macroscopic interaction coefficients satisfy the gauge-fixing conditions $V_{ij}^{r}=V_{ji}^{r}$ and $V_{ij}^{nr}=-V_{ji}^{nr}$. Taking the limit $\Omega\to\infty$ yields the deterministic limit, $\bar{V}_{ij}^{r}=\lim_{\Omega\to\infty}\beta\Omega v_{ij}^{r}$ and $\bar{V}_{ij}^{nr}=\lim_{\Omega\to\infty}\beta\Omega v_{ij}^{nr}$. Throughout this paper, we focus on the macroscopic coarse-grained description, though the system-specific scale $\Omega$ should be imposed depending on the physical context. In particular, when the mean number of particles per lattice site is small, a mesoscopic description is a better alternative for coarse-grained physical analysis.
@@
Occupancy noise Poissonian eq. 25 [87] Gaussian corrections around the mean field Vanishes, recovering the mean-field limit
Transition noise Poissonian and $O(1)$ [89, 90] Gaussian and $O(1/\Omega)$ Vanishes
Table 4: The summary of the impact and importance of fluctuations across different coarse-grained descriptions.
Table 4 summarizes the implications of fluctuations at different observation scales: the mesoscopic, macroscopic, and deterministic scales. The relevant coarse-grained states are the number of particles at a lattice site, the particle density at a lattice point, and the particle density at a lattice point, respectively. The suitable physical models corresponding to these scales are the Lattice Gas Models (LGM), Macroscopic Fluctuation Theory (MFT), and Chemical Reaction Networks (CRN). CRNs are a special case where $\Omega=\mathcal{V}$. While structurally similar, these descriptions differ significantly in other physical aspects, particularly concerning the nature of the noise effects.
Noise plays a key role through two mechanisms: particle occupancy and the transitions between them. Equation 25 highlights that Poissonian occupancy noise renormalizes the mesoscopic $V_{ij}$, leading to a nonlinear dependence on $v_{ij}$. It is crucial for correctly predicting the microscopic phase diagram using coarse-grained mesoscopic EOMs [87]. Moreover, the mean EPR correctly incorporates the microscopic noise effects of occupancy. $(\dagger)$ In comparison, the Gaussian/Langevin approximation eq. 29 of the mesoscopic Poissonian transition noise is sufficient due to the van Kampen closure approximation and the correct identification of the fluctuation-response relation, as discussed in the subsequent section. The non-quadratic dissipation function leads to a non-quadratic Hamilton-Jacobi equation [173, 174, 175, 176, 177], and eq. 29 incorporates this mesoscopic effect despite the Gaussian/Langevin formulation. A more systematic analysis of mesoscopic Poissonian transition fluctuations is detailed in [89, 90]. Importantly, we have not obtained the macroscopic EPR using the Langevin equation, which avoids the underestimation of the macroscopic EPR near equilibrium [89, 90]. In particular, the correct identification of effective transition affinities using the mean and fluctuations of the transition currents resolves this issue [89, 90].
@@ 4 Thermodynamic relations
Here, we outline different thermodynamic relations using the macroscopic description. However, one could also use the microscopic description, which is more directly relatable in Stochastic Thermodynamics. We deliberately utilize the macroscopic description to extend and exhibit the applicability of Stochastic Thermodynamics to interacting many-body systems.
@@ 4.1 Non-reciprocal Onsager relations
In this section, we demonstrate the non-reciprocal Onsager relation [178, 179]. Assuming vanishing external driving, $\vec{F}_{i}^{sp}=0$ and $F_{\gamma\gamma^{\prime}}^{ch}=0$, the $J^{\vec{\mathcal{D}}}_{i}$ in eq. 30 reads
$\vec{J}^{\vec{\mathcal{D}}}_{i}=-\sum_{j}\left(D_{ij}^{r}\left(\{\rho\}\right)+D_{ij}^{nr}\left(\{\rho\}\right)\right)\nabla^{\vec{\mathcal{D}}}\rho_{j},$ (44)
@@
where $D_{ij}^{r}(\{\rho\})=D_{i}^{\vec{{\mathcal{D}}}}(\{\rho\})(\beta V_{ij}^{r}+\delta_{ij}/\rho_{i})$ and $D_{ij}^{nr}(\{\rho\})=\beta D_{i}(\{\rho\})V_{ij}^{nr}$. The term $D_{ij}^{r}(\{\rho\})$ satisfies Onsager’s reciprocal relation [178, 179], since $V_{ij}^{r}={\partial^{2}E}/{\partial\rho_{j}\partial\rho_{i}}={\partial^{2}E}/{\partial\rho_{i}\partial\rho_{j}}$. In contrast, $D_{ij}^{nr}(\{\rho\})$ satisfies Onsager’s anti-reciprocal relation, due to $V_{ij}^{nr}=-V_{ji}^{nr}$, or equivalently $\partial F_{i}^{nr}/\partial\rho_{j}=-\partial F_{j}^{nr}/\partial\rho_{i}$. We introduce the mobility, entropic, symmetric, and anti-symmetric interaction matrices $\mathbb{D}$, $\mathbb{S}$, $\mathbb{D}^{r}$, and $\mathbb{D}^{nr}$, with $ij^{th}$ element of the matrices being $\delta_{ij}D_{i}(\{\rho\})$, $\delta_{ij}/\rho_{i}$, $D_{ij}^{r}$, and $D_{ij}^{nr}$, respectively, and let $\boldsymbol{\nabla\rho}$ denote the column vector of gradients.
[…]
@@
Using $\boldsymbol{\rho}=\boldsymbol{\rho}^{ss}+\delta\boldsymbol{\rho}$ and/or $\mathbf{J}^{\vec{\mathcal{D}}}=\mathbf{J}^{\vec{\mathcal{D}}}_{ss}+\delta\mathbf{J}^{\vec{\mathcal{D}}}$, and exploiting the symmetry of fluctuations around the steady-state profile $\boldsymbol{\rho}^{ss}$ or $\mathbf{J}^{\vec{\mathcal{D}}}_{ss}$, one can further decompose $\langle\dot{\Sigma}^{\mathcal{D}}\rangle$ into nonadiabatic and housekeeping contributions. Similarly, $\langle\dot{\Sigma}^{\mathcal{R}}\rangle$ can be incorporated by linearizing the reactive currents around equilibrium $\boldsymbol{\rho}^{eq}$ ($\mathbf{J}^{\vec{\mathcal{R}}}_{eq}$) or steady-state $\boldsymbol{\rho}^{ss}$ ($\mathbf{J}^{\vec{\mathcal{R}}}_{ss}$), requiring identification of analogous symmetric and anti-symmetric couplings $\mathbb{D}^{r}$ and $\mathbb{D}^{nr}$ in the discrete transition space [180].